The following intial rate data were obtained by adding 2.5 μg of a purified enzyme having molecular weight of 42,300 Daltons into 100 mL of water with the indicated concentrations of substrate C) So (mmol/L) 1.00 2.00 4.00 8.00 (HmoVLs) 4.50 7.90 11.10 15.20 Calculate the value of kCAT. KM and the pseudo first order rate constant.
Answer #1

Michaelis-menten equation:

Vo=Vmax [S]/Km+[S] where [S] =substrate concentration,Vmax=maximum velocity of enzyme-substrate reaction

Km=michaelis-menten constant,vo=reaction velocity

this can be rearranged to give : (1/Vo)=(Km/Vmax)*1/[S] +1/Vmax

A plot between 1/[S] vs 1/Vo is called lineweaver-Burk plot is linear with slope=Km/Vmax and y-intercept=1/Vmax

Also,Vmax=kcat*[E]

Vo=kcat*[E] [S]/Km+[S]

For high [S], km<<<[S]

So,Vo=kcat*[E] [S]/[S] =kcat [E]

thus kcat =pseudo first order rate constant as substrate is in large excess.

Data:

1/[So] (L/mmol)                       1/v(Ls/umol)

1.00                                            0.222

0.5                                               0.126

0.25                                             0.09

0.125                                           0.066

slope=km/Vmax=0.177 (Ls/umol) *(mmol/L)=0.177 *1000 (s)

intercept=1/Vmax=0.043 Ls/umol

Vmax=1/0.043=23.256 umol/Ls

km=0.177*1000 s *23.256 *10^-6 mol/Ls=4.116 *10^-3 mol/L=4.116 mmol/L

Also [E]=mol/volume

mol of enzyme=2.5ug/ 42300 g/mol=0.0000543 umol [1dalton=1g/mol]

[E]=0.0000543 umol/0.1L=0.000543 umol/L

So kcat=Vmax/[E]=(23.256 umol/Ls)/(0.000543 umol/L)=42828.729 s^-1

kcat=42828.729 s^-1

Spreadsheets Sign in Homeet Pae Layout Formulas Data Review View Special Features Cut Paste, 自Copy Format B 1 u B-隱 스·4 Merge